# A Formal Analysis of Algorithms for Matroids and Greedoids

### v1

@@ 4.1 Augmentation
For a matroid $(E,\mathcal{F})$ and $X\in\mathcal{F}$, if (1) $x_{1},...,x_{s}\in X$, (2) $y_{1},...,y_{s}\in E\setminus X$, (3) $x_{k}\in\mathcal{C}(X,y_{k})$, for all $1\leq k\leq s$ and (4) $x_{j}\not\in\mathcal{C}(X,y_{k})$, for all $1\leq j<k\leq s$. Then, $(X\setminus\{x_{1},...,x_{s}\}\cup\{y_{1},...,y_{s}\})\in\mathcal{F}$.
@@
The formalisation of the above lemma contains an inductive proof on $s$ within the context of the matroid locale matroid. The $x_{i}$s and $y_{i}$s are formalised as a list of pairs.
From now on, we assume two matroids $(E,\mathcal{F}_{1})$ and $(E,\mathcal{F}_{2})$ over the same ground set $E$, or formally, we work in the double_matroid locale (Listing LABEL:isabelle:inter_opt_def). For a set $X\in\mathcal{F}_{1}\cap\mathcal{F}_{2}$, we define an auxiliary graph $G_{X}$ with vertices $E$ and edges $A_{X,1}\cup A_{X,2}$ where $A_{X,1}=\{(x,y)\;.\;y\in E\setminus X\wedge x\in\mathcal{C}_{1}(X,y)\setminus\{y\}\}$ and $A_{X,2}=\{(y,x)\;.\;y\in E\setminus X\wedge x\in\mathcal{C}_{2}(X,y)\setminus\{y\}\}$. $G_{X}$ is obviously bipartite between $X$ and $E\setminus X$. We define two sets $S_{X}=\{y.\;y\in E\setminus X\wedge X\cup\{y\}\in\mathcal{F}_{1}\}$ and $T_{X}=\{y.\;y\in E\setminus X\wedge X\cup\{y\}\in\mathcal{F}_{2}\}$. A path between a vertex from $S_{X}$ and another from $T_{X}$ indicates an alternating sequence of insertion and deletion. Due to bipartiteness, the length will be odd allowing for an augmentation. We make this precise as follows.
Lemma 7 (Augmentation Lemma).
@@
The pairs of $x$s and $y$s are formalised as a zipped list. As indexing starts at 00 in Isabelle/HOL, the indices are often off by $1$. The statement is in Listing LABEL:isabelle:inter_opt_def.
@@
Theorem 8 (Optimality Criterion).
$X$ is a set of maximum cardinality in $\mathcal{F}_{1}\cap\mathcal{F}_{2}$ iff the $G_{X}$ does not contain a path from some $s\in S_{X}$ to some $t\in T_{X}$ (henceforth, an $S_{X}-T_{X}path)$. Edmonds’ max-min-equality [12] is a corollary: $\max\{|X|.\;X\in\mathcal{F}_{1}\cap\mathcal{F}_{2}\}=\min\{r_{1}(Q)+r_{2}(E\setminus Q).\;Q\subseteq E\}$.
@@ 4.2 Intersection Algorithm
We exploit the previous subsection’s results for an algorithm [6, 24] that iteratively applies augmenting paths to solve maximum cardinality matroid intersection. The main invariant of the algorithm MaxMatroidIntersection is $X\in\mathcal{F}_{1}\cap\mathcal{F}_{2}$ for the current solution $X$. Algorithm 2 contains the pseudocode. In our main reference [21], the circuits $\mathcal{C}_{1/2}(X,y)$, which are minimal dependent sets, are computed as explicit sets by iterating over all $x\in X\cup\{y\}$ and taking those where $X\cup\{y\}\setminus\{x\}$ is independent. $G_{X}$ is then computed by using the definition from above which requires further iterations over the circuits. Our pseudocode does not need the circuits as an intermediate step and adds an edge $(x,y)$ or $(y,x)$ if $X\cup\{y\}$ is dependent or $X\setminus\{x\}\cup\{y\}$ is independent, respectively.
@@ 1 Initialise $X\leftarrow\emptyset$;
MaxMatroidIntersection is again formalised within a locale that assumes the necessary subprocedures, most notably a function find_path for path selection. This will take the graph $G_{X}$ as an adjacency map. We also have two weak independence oracles orcl1 and orcl2, one for each matroid. We assume a Haskell-style outer_fold and inner_fold to simulate for-loops over sets in the functional language of Isabelle/HOL. We furthermore have a function to complement sets for the term $E\setminus X$.
Listing LABEL:isabelle:intersec_algo shows the formalisation of the for-loops. treat1 and treat2 simulate the for-loops in Line 6 and Line 8, respectively. The outer for-loop extending from Lines 4 to 8 is realised by compute_graph. We also have a simple function augment performing the augmentation, i.e. alternating deletion and insertion to a set. The function matroid_intersection (while loop) calls the graph computation, gives $A_{X,1}\cup A_{X,2}$, $S_{X}$ and $T_{X}$ to the path selection function and terminates or calls augmentation according to the result of path selection. It maps the solution $X$ before entering the loop to the $X$ after finishing the loop. We use a record with a single variable for the current solution.
Again, we use invariants for loop verification. We have independence of $X$ as the major invariant and data structure well-formedness invariants as minor ones. As in the case of BestInGreedy, we use the approach of proving single-step preservation for execution paths that is combined by an induction. The first lemma in Listing LABEL:isabelle:intersec_invar says that by executing one step (a) $X$ remains in $\mathcal{F}_{1}\cap\mathcal{F}_{2}$, (b) $X$ continues to satisfy its data structure invariant, (c) $X\subseteq E$, (d) $X$’s cardinality increases by $1$ and (e) neither $S_{X}$ nor $T_{X}$ are empty. The Listing also contains the lemma certifying optimality if the terminating branch of matroid_intersection is taken. The second lemma together with (a), (b) and (c) can be combined by the induction principle into correctness for general inputs $X$. When (d) and (e) are used, there is termination. Finally, one can prove total correctness when matroid_intersection is called on initial_state.
@@
Running Time and Circuit Oracles The number of iterations to build $G_{X}$ is $\mathcal{O}(|E|^{2})$. Each matroid comes with an oracle which is assumed to be $\iota_{1}$ or $\iota_{2}$, respectively. This results in $\mathcal{O}(|E|^{3}\cdot(\iota_{1}+\iota_{2}))$ overall running time. We might have to check $|X|\cdot(|E|-|X|)$ pairs of vertices. If - depending on the problem - the size of the circuits $\sigma_{1}$ and $\sigma_{2}$ is assumed small, e.g. $\mathcal{O}(\log|E|)$, most of the inner for-loops’ iterations were wasted. If the circuits can be computed in time $\kappa_{1},\kappa_{2}\in o(|E|)$, the time would be the more efficient $\mathcal{O}(|E|^{2}\cdot(\iota_{1}+\iota_{2}+\kappa_{1}+\kappa_{2}+\sigma_{1}+\sigma_{2}))$.
@@
Selecting an $S_{X}$-$T_{X}$ Path There is a verified implementation of Breadth-First Search (BFS) that works on adjacency maps by Abdulaziz  [29]. It takes a set of sources and builds another adjacency map modelling a DAG of those edges that are actually explored by the search. If a vertex $v$ is reachable from a source $s$, there is a path $p$ in this DAG leading from a source $s^{\prime}$ to $v$ such that $p$ is a shortest path in the actual graph leading from $s^{\prime}$ to $v$. find_path runs BFS for the sources $S_{X}$ and uses a DFS to obtain a path in the DAG and, hence, a shortest $S_{X}$-$T_{X}$ path. For a concrete intersection problem and together with appropriate oracles, find_path can be used to instantiate the locale of the intersection algorithm.
Bipartite Matching An example for matroid intersection is maximum cardinality bipartite matching. Consider a bipartite graph with edges $E$ between two disjoint sets of vertices $L$ and $R$. For $e\in E$, we call the endpoint in $L$ the left and the one in $R$ the right endpoint. We say that $M\subseteq E$ is independent w.r.t. $L$ iff no edges in $M$ share a left endpoint. Independence w.r.t. $R$ is defined analogously. Valid matchings are sets of vertex-disjoint edges in $E$ and are thus exactly those $M$ that are independent w.r.t. both $L$ and $R$. The two independence predicates satisfy the matroid axioms making this a matroid intersection problem. Circuits w.r.t. $L$ or $R$ would be edges $e,d\in E$ sharing a left or right endpoint, respectively.
@@ 5 Discussion
We presented a formal analysis of algorithms to solve two optimisation problems for matroids: (a) computing the largest (in terms of accumulated weight) independent set in a given matroid and (b) computing the largest (in terms of cardinality) set that is independent w.r.t. two matroids. The two algorithms are of great importance to practitioners and are implemented in a number of computer algebra systems, e.g. in Sage [1] and Macaulay [2]. Additionally, we also briefly (due to space constraints) presented a formalisation of the analysis of an algorithm to solve an optimisation problem for greedoids, which are a generalisation of matroids. In addition to formalising significant parts of the theory of matroids, greedoids, and the analysis of the algorithms, we showed that our approach can also be used to obtain practical executable verified implementations of important graph algorithms. This demonstrates the potential practical role of matroids and greedoids to design well-factored formal mathematical libraries whereby proofs of multiple algorithms are done in one go that captures their mathematical essence, and that is later instantiated for different concrete examples. The formalisation overall totalled 17.4K lines of proof script, with 11K lines dedicated to the theory of matroids and greedoids, 2.9K lines connecting graph theoretic problems to matroids and/or greedoids, and 3.5K lines on defining and verifying Kruskal’s, Prim’s, and the bi-partite matching algorithm. Our formalisation builds on and is part of an ongoing effort to formalise combinatorial optimisation in Isabelle/HOL[4].
@@
Related work Most relevant to us is the formalisation of matroids in Isabelle/HOL by Keinholz [16]. Haslbeck et al. [14] formalised a classic proof of the correctness of the Best-In-Greedy Algorithm and obtained a verified, imperative implementation of Kruskal’s algorithm. They used a weak oracle based on a verified imperative union-find implementation by Löwenberg [25] to store the tree’s connected components leading to their implementation having the best possible running time of $\mathcal{O}(m\alpha(n))$, where $\alpha$ is the inverse Ackermann function. In Lean, there is an ongoing effort to formalise matroid theory [28], including more advanced results about e.g. minors, isomorphism, Tutte’s excluded minor theorem for finitary binary matroids, or Edmonds’ Intersection Theorem (Lemma 5), but (as of the time of writing) there seems to be no verified practical algorithms in that library. Matroids were also formalised in Coq [26, 9] and Mizar [8], however, both do not go very far. Set systems were formalised in Isabelle/HOL by Edmonds and Paulson [10], who did that in the context of verifying combinatorial designs. Additionally, there are formalisations of matching algorithms [3, 5] and Prim’s algorithm [23], in Isabelle/HOL, and of Dijkstra, Kruskal, and Prim [27], in Coq, that are not based on matroids or greedoids, where all reasoning is done at the graph-theoretic level. Readers interested in comparing the style of formal reasoning when using matroids or greedoids, which is algebraic, to the one performed directly on graphs, which is more intuitive but combinatorial, should refer to those formalisations.
@@
Future work There is one ’main’ missing algorithm from our library which would be the most imminent piece of future work: the weighted version of the matroid intersection problem, whereby, instead of finding the largest cardinality set that is independent w.r.t. two matroids, one is to find the largest in terms of the weight of such a set, for a modular cost function.
@@
Routine functions that check whether a matroid indeed satisfies the assumptions of a matroid, compute the dual of a matroid, etc., which are implemented in most computer algebra packages, are missing from our library and are an important next step for our work. Naive implementations of those functions would have a worst-case exponential running time, so utilising sophisticated enumeration algorithms [17, 18] would help with making that more practical.

### v2

@@ 4.1 Augmentation
For a matroid $(E,\mathcal{F})$ and $X\in\mathcal{F}$, assume (1) $x_{1},...,x_{s}\in X$, (2) $y_{1},...,y_{s}\in E\setminus X$, (3) $x_{k}\in\mathcal{C}(X,y_{k})$, for all $1\leq k\leq s$, and (4) $x_{j}\not\in\mathcal{C}(X,y_{k})$ for all $1\leq j<k\leq s$. If (1)-(4) hold, then $(X\setminus\{x_{1},...,x_{s}\}\cup\{y_{1},...,y_{s}\})\in\mathcal{F}$.
@@
The formalisation of the above lemma contains an inductive proof on $s$ within the context of the locale matroid. The $x_{i}$s and $y_{i}$s are formalised as a list of pairs.
From now on, we assume two matroids $(E,\mathcal{F}_{1})$ and $(E,\mathcal{F}_{2})$ over the same ground set $E$, or formally, we work in the double_matroid locale (Listing LABEL:isabelle:inter_opt_def). Following Korte and Vygen [21], for a set $X\in\mathcal{F}_{1}\cap\mathcal{F}_{2}$, we define an auxiliary graph $G_{X}$ with vertices $E$ and edges $A_{X,1}\cup A_{X,2}$ where $A_{X,1}=\{(x,y)\;.\;y\in E\setminus X\wedge x\in\mathcal{C}_{1}(X,y)\setminus\{y\}\}$ and $A_{X,2}=\{(y,x)\;.\;y\in E\setminus X\wedge x\in\mathcal{C}_{2}(X,y)\setminus\{y\}\}$. $G_{X}$ is obviously bipartite between $X$ and $E\setminus X$. We define two sets $S_{X}=\{y.\;y\in E\setminus X\wedge X\cup\{y\}\in\mathcal{F}_{1}\}$ and $T_{X}=\{y.\;y\in E\setminus X\wedge X\cup\{y\}\in\mathcal{F}_{2}\}$. A path between a vertex from $S_{X}$ and another from $T_{X}$ indicates an alternating sequence of insertion and deletion. Due to bipartiteness, the length will be odd allowing for an augmentation. We make this precise as follows.
Lemma 7 (Augmentation Lemma [21]).
@@
The statement is in Listing LABEL:isabelle:inter_opt_def. We show that the edges of paths used by Lemma 7 alternate between $A_{1}$ and $A_{2}$ using an existing library on alternating lists for matchings [4]. From this we can easily deduce (1)-(3) to apply Lemma 6. The re-applicability of matching theory is no coincidence since bipartite matching is an instance of matroid intersection as we show later.
@@
Theorem 8 (Optimality Criterion [21]).
$X$ is a set of maximum cardinality in $\mathcal{F}_{1}\cap\mathcal{F}_{2}$ iff $G_{X}$ does not contain a path from some $s\in S_{X}$ to some $t\in T_{X}$ (henceforth, an $S_{X}$-$T_{X}$ path). Edmonds’ max-min-equality [13] is a corollary: $\max\{|X|.\;X\in\mathcal{F}_{1}\cap\mathcal{F}_{2}\}=\min\{r_{1}(Q)+r_{2}(E\setminus Q).\;Q\subseteq E\}$.
@@ 4.2 Intersection Algorithm
We exploit the previous subsection’s results for an algorithm [7, 24] that iteratively applies augmenting paths to solve maximum cardinality matroid intersection. The main invariant of the algorithm MaxMatroidIntersection is $X\in\mathcal{F}_{1}\cap\mathcal{F}_{2}$ for the current solution $X$ (preservation proof by Lemma 7). Algorithm 2 shows the pseudocode. In our main reference [21], the circuits $\mathcal{C}_{1/2}(X,y)$, i.e. minimal dependent sets, are computed as explicit sets by iterating over all $x\in X\cup\{y\}$ and taking those where $X\cup\{y\}\setminus\{x\}$ is independent. $G_{X}$ is then computed by using the definition from above which requires further iterations over the circuits. Our pseudocode does not need the circuits as an intermediate step and adds an edge $(x,y)$ or $(y,x)$ if $X\cup\{y\}$ is dependent or $X\setminus\{x\}\cup\{y\}$ is independent, respectively.
@@ 1 Initialise $X\leftarrow\emptyset$;
MaxMatroidIntersection is again formalised within a locale that assumes the necessary subprocedures, most notably a function find_path for path selection. This will take the graph $G_{X}$ as an adjacency map. We also have two weak independence oracles orcl1 and orcl2, one for each matroid. We assume a Haskell-style outer_fold and inner_fold to simulate for-loops over sets in the functional language of Isabelle/HOL. We furthermore assume a function to compute set complements as required in Line 4 of Algorithm 2.
Listing LABEL:isabelle:intersec_algo shows the formalisation of the for-loops. treat1 and treat2 simulate Line 6 and Line 8, respectively. The outer for-loop between Lines 4 to 8 is realised by compute_graph. We also have a simple function augment performing the augmentation, i.e. alternating deletion and insertion to a set. The function matroid_intersection (while loop) calls the graph computation, gives $A_{X,1}\cup A_{X,2}$, $S_{X}$ and $T_{X}$ to the path selection function and terminates or calls augmentation according to the result of path selection. matroid_intersection maps the solution $X$ before entering the loop (intended for Line 1 of Algorithm 2) to the $X$ after finishing the loop. We use a record with a single variable for the current solution.
Again, we use invariants for loop verification. Independence of $X$ is the major invariant and data structure well-formedness invariants are minor ones. As for BestInGreedy, we prove single-step preservation for execution paths which is then lifted to preservation for the whole function by a computational induction. The first lemma in Listing LABEL:isabelle:intersec_invar says that by executing one step (a) $X$ remains in $\mathcal{F}_{1}\cap\mathcal{F}_{2}$ (by Lemma 7), (b) $X$ continues to satisfy its data structure invariant, (c) $X\subseteq E$, (d) $X$’s cardinality increases by $1$ (by Lemma 7) and (e) neither $S_{X}$ nor $T_{X}$ are empty. The Listing also contains the lemma certifying optimality if the terminating branch of matroid_intersection is taken (by Theorem 8). The second lemma together with (a), (b) and (c) can be combined by the induction principle into correctness for general inputs $X$. When (d) and (e) are used, there is termination. Total correctness is proven when matroid_intersection is called on initial_state.
@@
Running Time and Circuit Oracles
The number of iterations to build $G_{X}$ is $\mathcal{O}(|E|^{2})$. Each matroid comes with an oracle which is assumed to be $\iota_{1}$ or $\iota_{2}$, respectively. This results in $\mathcal{O}(|E|^{3}\cdot(\iota_{1}+\iota_{2}))$ overall running time. We might have to check $|X|\cdot(|E|-|X|)$ pairs of vertices. If - depending on the problem - the size of the circuits $\sigma_{1}$ and $\sigma_{2}$ is assumed to be small, e.g. $\mathcal{O}(\log|E|)$ or even $\mathcal{O}(1)$, most of the inner for-loops’ iterations were wasted. If the circuits can be computed in time $\kappa_{1},\kappa_{2}\in o(|E|)$, the time would be the more efficient $\mathcal{O}(|E|^{2}\cdot(\iota_{1}+\iota_{2}+\kappa_{1}+\kappa_{2}+\sigma_{1}+\sigma_{2}))$.
@@
Selecting an $S_{X}$-$T_{X}$ Path
There is a verified implementation of Breadth-First Search (BFS) that works on adjacency maps by Abdulaziz [5]. It takes a set of sources and builds another adjacency map modelling a DAG of those edges that are actually explored by the search. If a vertex $v$ is reachable from a source $s$, there is a path $p$ in this DAG leading from a source $s^{\prime}$ to $v$ such that $p$ is a shortest path in the actual graph leading from $s^{\prime}$ to $v$. find_path runs BFS for the sources $S_{X}$ and uses a DFS to obtain a path in the DAG and, hence, a shortest $S_{X}$-$T_{X}$ path. For a concrete intersection problem and together with appropriate oracles, find_path can be used to instantiate the locale of the intersection algorithm.
Bipartite Matching
An example for matroid intersection is maximum cardinality bipartite matching. Consider a bipartite graph with edges $E$ between two disjoint sets of vertices $L$ and $R$. For $e\in E$, we call the endpoint in $L$ the left and the one in $R$ the right endpoint. We say that $M\subseteq E$ is independent w.r.t. $L$ iff no edges in $M$ share a left endpoint. Independence w.r.t. $R$ is defined analogously. Valid matchings are sets of vertex-disjoint edges in $E$ and are thus exactly those $M$ that are independent w.r.t. both $L$ and $R$. The two independence predicates satisfy the matroid axioms making this a matroid intersection problem. Circuits w.r.t. $L$ or $R$ would be edges $e,d\in E$ sharing a left or right endpoint, respectively.
@@ 5 Discussion
We presented a formal analysis of algorithms to solve two optimisation problems for matroids: (a) computing the largest (in terms of accumulated weight) independent set in a given matroid and (b) computing the largest (in terms of cardinality) set that is independent w.r.t. two matroids. The two algorithms are of great importance to practitioners and are implemented in a number of computer algebra systems, e.g. in Sage [1] and Macaulay [2]. Additionally, we also briefly (due to space constraints) presented a formalisation of the analysis of an algorithm to solve an optimisation problem for greedoids, which are a generalisation of matroids. In addition to formalising significant parts of the theory of matroids, greedoids, and the analysis of the algorithms, we showed that our approach can also be used to obtain practical executable verified implementations of important graph algorithms. This demonstrates the potential practical role of matroids and greedoids to design well-factored formal mathematical libraries whereby proofs of multiple algorithms are done in one go that captures their mathematical essence, and that is later instantiated for different concrete examples. The formalisation overall totalled 17.4K lines of proof script, with 11K lines dedicated to the theory of matroids and greedoids, 2.9K lines connecting graph theoretic problems to matroids and/or greedoids, and 3.5K lines on defining and verifying Kruskal’s, Prim’s, and the bi-partite matching algorithm. Our formalisation is available at https://doi.org/10.5281/zenodo.15758230; it builds on and is part of an ongoing effort to formalise combinatorial optimisation in Isabelle/HOL [3].
Related Work
@@
Most relevant to us is the formalisation of matroids in Isabelle/HOL by Keinholz [17]. Haslbeck et al. [15] formalised a classic proof of the correctness of the Best-In-Greedy Algorithm and obtained a verified, imperative implementation of Kruskal’s algorithm. They used a weak oracle based on a verified imperative union-find implementation by Löwenberg [25] to store the tree’s connected components leading to their implementation having the best possible running time of $\mathcal{O}(m\alpha(n))$, where $\alpha$ is the inverse Ackermann function. Löwenberg also gives a functional version that we could use for our work, which would, however, not pay off in terms of running time.
In Lean, there is an ongoing effort by Nelson et al. to formalise matroid theory [28], including more advanced results about e.g. minors, isomorphism, Tutte’s excluded minor theorem for finitary binary matroids, or Edmonds’ Intersection Theorem (Lemma 5), but (as of the time of writing) there seems to be no verified practical algorithms in that library.
Matroids were also formalised in Coq [26, 10] and Mizar [9], however, both do not go very far. Set systems were formalised in Isabelle/HOL by Edmonds and Paulson [11], who did that in the context of verifying combinatorial designs.
Additionally, there are formalisations of matching algorithms [4, 6] and Prim’s algorithm [23], in Isabelle/HOL, and of Dijkstra, Kruskal, and Prim [27], in Coq, that are not based on matroids or greedoids, where all reasoning is done at the graph-theoretic level. The Isabelle/HOL formalisation of Prim’s algorithm [27] uses a priority queue to find edges to extend the spanning tree. We cannot use this more efficient (and standard) way due to conditions required to prove Theorem 4. Readers interested in comparing the style of formal reasoning when using matroids or greedoids, which is algebraic, to the one performed directly on graphs, which is more intuitive but combinatorial, should refer to those formalisations.
Future Work
@@
There is one ’main’ missing algorithm from our library which would be the most imminent piece of future work: the weighted version of the matroid intersection problem, whereby, instead of finding the largest cardinality set that is independent w.r.t. two matroids, one is to find the largest in terms of the weight of such a set, for a modular cost function.
@@
Routine functions that check whether a matroid indeed satisfies the assumptions of a matroid, compute the dual of a matroid, etc., which are implemented in most computer algebra packages, are missing from our library and are an important next step for our work. Naive implementations of those functions have a worst-case exponential running time, so utilising sophisticated enumeration algorithms [18] would help with making that more practical.